# FlowMap: High-Quality Camera Poses, Intrinsics, and Depth via Gradient Descent

### v1

@@ 1 Introduction
Through extensive ablation studies, we show that each of FlowMap’s design choices is necessary. On popular, real-world novel view synthesis datasets (Tanks & Temples, Mip-NeRF 360, CO3D, and LLFF), we demonstrate that FlowMap enables photo-realistic novel view synthesis up to full $360^{\circ}$ trajectories using Gaussian Splatting [28]. Gaussian Splats obtained from FlowMap reconstructions far outperform the state-of-the-art gradient-based bundle-adjustment method, NoPeNeRF [2], and those obtained using the SLAM algorithm DROID-SLAM [65], even though both baselines require ground-truth intrinsics. Gaussian Splats obtained from FlowMap are on par with those obtained from COLMAP [56], even though FlowMap only leverages gradient descent, is fully differentiable, and represents a complete departure from conventional SfM techniques.
@@ 2 Related Work
Prior work has attempted to embed the full SLAM pipeline into a deep learning framework [62, 14, 80, 13, 67, 35, 63, 70, 3], usually by training black-box neural networks to directly output camera poses. However, these methods are constrained to short videos of 5 to 10 frames and are not competitive with conventional SLAM and SfM for real-world 3D reconstruction. Bowen et al.[4] elegantly leverage optical flow supervision for self-supervised monocular depth prediction. More recently, DROID-SLAM [65] has yielded high-quality camera poses and depth. However, it requires known intrinsics, is trained fully supervised with ground-truth camera poses, and fails to approach COLMAP on in-the-wild performance and robustness. Concurrent work to FlowMap explores an end-to-differentiable, point-tracking-based SfM framework [68]. Key differences are that their method is fully supervised with camera poses, point clouds, and intrinsics; requires large-scale, multi-stage training; solves only for sparse depth; and is built around the philosophy of making each part of the conventional SfM pipeline differentiable. Our method is a complete departure from the conventional SfM pipeline—it does not require a training set of known intrinsics, ground-truth poses, or 3D points, and it provides quality gradients for dense depth, poses, and intrinsics. Critically, FlowMap is the first gradient-descent based method to rival the performance of conventional SfM on the novel view synthesis task.
@@
A recent line of work in bundle-adjusting radiance fields [33, 2, 78, 27, 25, 21, 74, 22, 76, 75, 72, 10, 9] attempts to jointly optimize unknown camera poses and radiance fields. Jeong et al. [25] additionally solve for camera intrinsic parameters. However, these methods only succeed when given forward-facing scenes or roughly correct pose initializations. More recent work incorporates optical flow and monocular depth priors [2, 40, 36] but requires known intrinsics and only works robustly on forward-facing scenes. Concurrent work [22] somewhat accelerates optimization compared to earlier NeRF-based approaches, but still requires intrinsics and yields significantly worse poses than COLMAP on 360-degree sequences. Further concurrent work proposes real-time SLAM via gradient descent on 3D Gaussians [39], but requires known intrinsics and does not show robustness on a variety of real-world scenes. In contrast, our method is robust and easily succeeds on object-centric scenes where the camera trajectory covers a full $360^{\circ}$ of rotation, yielding photo-realistic novel view synthesis when combined with Gaussian Splatting.
@@ 7.0.3 Pre-training Depth and Correspondence Networks (Q3)
8 Conclusion
@@
This work was supported by the National Science Foundation under Grant No. 2211259, by the Singapore DSTA under DST00OECI20300823 (New Representations for Vision and 3D Self-Supervised Learning for Label-Efficient Vision), by the Intelligence Advanced Research Projects Activity (IARPA) via Department of Interior/ Interior Business Center (DOI/IBC) under 140D0423C0075, by the Amazon Science Hub, and by IBM. The Toyota Research Institute also partially supported this work. The views and conclusions contained herein reflect the opinions and conclusions of its authors and no other entity.
@@
For our depth network, we use the lightweight CNN version of MiDaS [50], pretrained with the publicly available weights trained on relative-depth estimation.

### v2

@@ 1 Introduction
Through extensive ablation studies, we show that each of FlowMap’s design choices is necessary. On popular, real-world novel view synthesis datasets (Tanks & Temples, Mip-NeRF 360, CO3D, and LLFF), we demonstrate that FlowMap enables photo-realistic novel view synthesis up to full $360^{\circ}$ trajectories using Gaussian Splatting [29]. Gaussian Splats obtained from FlowMap reconstructions far outperform the state-of-the-art gradient-based bundle-adjustment method, NoPe-NeRF [2], and those obtained using the SLAM algorithm DROID-SLAM [67], even though both baselines require ground-truth intrinsics. Gaussian Splats obtained from FlowMap are on par with those obtained from COLMAP [58], even though FlowMap only leverages gradient descent, is fully differentiable, and represents a complete departure from conventional SfM techniques.
@@ 2 Related Work
Prior work has attempted to embed the full SLAM pipeline into a deep learning framework [64, 14, 84, 13, 69, 37, 65, 72, 3], usually by training black-box neural networks to directly output camera poses. However, these methods are constrained to short videos of 5 to 10 frames and are not competitive with conventional SLAM and SfM for real-world 3D reconstruction. Bowen et al.[4] elegantly leverage optical flow supervision for self-supervised monocular depth prediction. More recently, DROID-SLAM [67] has yielded high-quality camera poses and depth. However, it requires known intrinsics, is trained fully supervised with ground-truth camera poses, and fails to approach COLMAP on in-the-wild performance and robustness. Concurrent work to FlowMap explores an end-to-differentiable, point-tracking-based SfM framework [70]. Key differences are that their method is fully supervised with camera poses, point clouds, and intrinsics; requires large-scale, multi-stage training; solves only for sparse depth; and is built around the philosophy of making each part of the conventional SfM pipeline differentiable. Our method is a complete departure from the conventional SfM pipeline—it does not require a training set of known intrinsics, ground-truth poses, or 3D points, and it provides quality gradients for dense depth, poses, and intrinsics. Critically, FlowMap is the first gradient-descent based method to rival the performance of conventional SfM on the novel view synthesis task. Zhang et al. [82] and Kopf ef al. [32] demonstrate gradient-descent based optimization of camera parameters with a similar flow-based reprojection supervision, with a focus on dynamic scenes. However, these methods optimize camera parameters as free variables and depend on pre-trained monocular depth estimators. In constrast, our feed-forward parameterization uniquely enables gradients for large-scale training and we demonstrate that our gradients can be used to train a depth estimator.
@@
A recent line of work in bundle-adjusting radiance fields [35, 2, 81, 28, 26, 21, 76, 22, 79, 77, 74, 10, 9] attempts to jointly optimize unknown camera poses and radiance fields. Several of these methods  [26, 78, 24] additionally solve for camera intrinsic parameters. However, these methods only succeed when given forward-facing scenes or roughly correct pose initializations. More recent work incorporates optical flow and monocular depth priors [2, 42, 38] but requires known intrinsics and only works robustly on forward-facing scenes. Concurrent work [22] accelerates optimization compared to earlier NeRF-based approaches. Unlike ours, this approach requires known intrinsics and a pre-trained monocular depth estimator, and minimizes a volume-rendering-based photometric loss instead of the proposed correspondence-based approach. Further concurrent work proposes real-time SLAM via gradient descent on 3D Gaussians [41], but requires known intrinsics and does not show robustness on a variety of real-world scenes. In contrast, our method is robust and easily succeeds on object-centric scenes where the camera trajectory covers a full $360^{\circ}$ of rotation, yielding photo-realistic novel view synthesis when combined with Gaussian Splatting.
@@ 7.0.3 Pre-training Depth and Correspondence Networks (Q3)
8 Discussion
8.0.1 Limitations.
FlowMap has several limitations that suggest exciting directions for future work. First, FlowMap requires off-the-shelf correspondences from optical flow and point tracking methods. An exciting direction is to remove the dependence on correspondence altogether or to jointly learn correspondence extraction. Second, we mainly analyze FlowMap in the setting of per-scene optimization, where our results demonstrate that the gradients provided by FlowMap’s formulation are robustly lead to high-quality depth and camera parameters. It is natural to attempt to use these gradients to train a feed-forward structure-from-motion method. Lastly, through its dependence on optical flow or point tracks, FlowMap can currently only process continuous video, in contrast to conventional SfM methods which can operate on unstructured image collections. This is a natural assumption for applications in embodied intelligence, navigation, and robotics, but limits applications in computer graphics. Leveraging unstructured correspondences, e.g. via [55], may be used to overcome this limitation.
8.0.2 Relationship to Conventional SfM.
Across many applications of conventional SfM, such as the reconstruction of large, non-continuous image collections, FlowMap cannot serve as a drop-in replacement, and we note that this is not our objective. Rather, we demonstrate that a self-supervised, end-to-end differentiable, feed-forward formulation that can naturally be integrated into neural network vision models surprisingly approaches COLMAP’s performance on the downstream task of novel view synthesis in the context of video data. Here, FlowMap has the potential to make camera pose and depth supervision unnecessary for 3D deep learning, paving the way for training on unannotated, internet-scale video data.
8.0.3 Conclusion.
@@ 8 Discussion
This work was supported by the National Science Foundation under Grant No. 2211259, by the Singapore DSTA under DST00OECI20300823 (New Representations for Vision and 3D Self-Supervised Learning for Label-Efficient Vision), by the Intelligence Advanced Research Projects Activity (IARPA) via Department of Interior/ Interior Business Center (DOI/IBC) under 140D0423C0075, by the Amazon Science Hub, and by IBM. The Toyota Research Institute also partially supported this work. The views and conclusions contained herein reflect the opinions and conclusions of its authors and no other entity. Vincent thanks his Mom for crocheting baby Yoda.
@@
For our depth network, we use the lightweight CNN version of MiDaS [52], pretrained with the publicly available weights trained on relative-depth estimation. We optimize the entire network weights during training.